OpenStudy (watermelon_talk):

Ari’s chickens eat their pellets at a constant rate throughout each day. What is the slope of the line that represents the ounces of pellets eaten? Ari’s Chicken Pellets Days 2 4 6 8 Ounces of Pellets remaining 39 33 27 21

OpenStudy (watermelon_talk):

@imqwerty

OpenStudy (karolynxx):

The general equation for a linear graph is y = ax + b a is the slope a is just the difference between each number, which is -6 as the numbers are decreasing by 6 each time so the slope is -6 but slope is actually -3 as the day gap is 2 so slope = -6/2 = -3